# Unlocking hidden biomolecular conformational landscapes in diffusion models at inference time

### v1

@@ 1 Introduction
Biology depends on molecular flexibility. Proteins, RNA, DNA, and other components of biological systems adopt dynamic 3D conformations, which are key to function [henzler-wildman_dynamic_2007, al-hashimi_rna_2008]. Understanding and modeling these dynamics informs both basic biology research and applications to medicine, agriculture, and other areas, and is therefore a major aim of both experimental and computational scientists. The statistical distribution of the atomic configurations of a molecular system in equilibrium at a given temperature is determined by their potential energies. Specifically, a state with lower potential energy is exponentially more probable than a higher potential energy state, as determined by the Boltzmann distribution [tuckerman_statistical_2023]. To sample this distribution, scientists have invested decades of effort in creating methods such as Monte Carlo samplers and molecular dynamics simulations (MD), but these methods are typically extremely slow and suffer from accuracy problems [tuckerman_statistical_2023, ormeno_convergence_2024, van_gunsteren_methods_2024].
More recently, machine learning models have revolutionized biomolecular structure prediction, and efforts are underway to extend these advances to train models that do not merely predict static structures but generate diverse conformations from the true Boltzmann distribution [jumper_highly_2021, abramson_accurate_2024, jing_alphafold_2023, wayment-steele_predicting_2024, lewis_scalable_2025]. As the field evolves, new questions arise about how best to take advantage of these models, access the information they contain, and target opportunities for further improvement. We ask the following question: Given a pretrained generative structure model, what strategy should be deployed at inference time to efficiently and accurately sample its conformational landscape?
In this work, we present an enhanced sampling algorithm, ConforMix, that can be applied to diffusion-based biomolecular structure prediction models such as the AlphaFold 3 family. ConforMix reveals hidden conformations and estimates free energies more efficiently than the default sampling from diffusion models. Our approach is based on the hypothesis that the probability distribution of generative models contains more information than is efficiently sampled. Importantly, even a generative model that perfectly sampled the true Boltzmann distribution—a major long-term goal in machine learning for biology—would, for many uses, still require enhanced sampling algorithms to fruitfully explore it. For instance, correct sampling of the tight-binding proteins barnase and barstar at standard state would generate an unbound structure with probability approximately $10^{-14}$ [wang_how_2004]. Directly drawing samples from the distribution would be an extremely inefficient way to estimate this probability, determine transition states along the binding pathway, or understand the effects of environmental factors in modulating this affinity [zhou_modeling_2013, trendelkamp-schroer_efficient_2016].
@@ 2 Related Work
Computational methods have been used to sample molecular conformations for decades. The well-known Metropolis-Hastings algorithm [metropolis_equation_1953] for Monte Carlo sampling was first used to simulate the statistical mechanics of particles in a box. Elastic network models represent a protein as low-resolution beads connected by springs and recover the principal motions that would occur in this representation, which often are representative of motions of the actual system [lopez-blanco_new_2016]. More sophisticated methods, such as coarse-grained or all-atom molecular dynamics simulation, model the physical interactions with progressively more detail and have been used to generate detailed probabilistic distributions for a number of proteins [lindorff-larsen_how_2011, mirarchi_mdcath_2024].
More recently, machine learning models such as AlphaFold 3 [abramson_accurate_2024] have revolutionized the problem of static protein structure prediction. AlphaFold 3 is the most prominent example of a class of diffusion-based structure prediction methods that excel at generating the single most likely protein structure for a given amino acid sequence. While AlphaFold 3 generates five samples per run, for well-ordered proteins these samples are often almost identical and are better understood as very similar predictions of the same static structure rather than a representation of conformational variety.
Using such models to model the distribution of conformations, however, is a task still in its infancy. The most well-studied approaches involve subsampling or otherwise perturbing the multiple sequence alignment (MSA) input [wayment-steele_predicting_2024, roney_state---art_2022, kalakoti_afsample2_2025, lee_large-scale_2025] or templates input [heo_multi-state_2022, del_alamo_sampling_2022, chiesa_modeling_2025] to AlphaFold and related models. These inputs specify, respectively, sequence co-evolutionary information and related 3D structures. Emphasizing different sequence or structural features in the input has empirically produced significant conformational diversity. Subsampling methods, however, suffer from fundamental limitations. First, subsampling necessarily involves reducing the information available to the model, which often produces poor-quality outputs. Second, molecular conformations are continuous, while subsampling of sequences or templates is a discrete operation. Third, the structure distribution accessed by subsampling methods is not well defined, and despite years of study there is no clear means of reconstructing actual probabilistic ensembles.
Beyond input-perturbation methods, other efforts have involved building new models to predict conformations [lewis_scalable_2025, fan_accurate_2024, jing_alphafold_2023]. Strategies include leveraging molecular dynamics simulations or experimental data, either during training or at runtime [liu_exendiff_2024, brown_approximating_2024, wang_protein_2024, lu_str2str_2024, levy_solving_2025]. While these approaches have shown promise, they do not yet accurately reproduce the thermodynamic ensembles of arbitrary proteins.
@@ 3.1 ConforMix
Diffusion models learn probability distributions, and the standard sampling procedures, which we refer to as default sampling, are designed to draw i.i.d. samples from this distribution. To explore further and more efficiently than is possible with default sampling, we employ conditional sampling, which is widely used with diffusion models to guide sampling toward particular regions of the learned probability distribution [dhariwal_diffusion_2021, ho_classifier-free_2022]. As we will see, conditional sampling offers clear benefits even without specialized prior knowledge about the protein of interest. While time-dependent classifier guidance and classifier-free guidance are commonly used in other diffusion contexts, these methods are not well suited to our needs because they require either additional training or special architectural choices, which are impractical for our needs.
Instead, to devise a purely inference-time method, we turn to Twisted Diffusion Sampling [wu_practical_2024], a form of sequential Monte Carlo (SMC), to guide conditional generation. SMC has advantages over more typical guidance methods, including (a) requiring only a time-independent guidance function, which enables us to easily describe conditions as a function of the fully denoised state of the system, and (b) providing an asymptotic guarantee of sampling the true target conditional distribution in the limit of many particles. Twisted diffusion sampling works in conjunction with an underlying diffusion model that learns to sample a target data distribution $q(x_{0}|s).$ Diffusion models define a forward noising process, operating over a sequence of $T$ steps, that constructs an extended distribution $q(x_{T:0}|s)=q(x_{T}|x_{T_{1}})\cdots q(x_{1}|x_{0})q(x_{0}|s).$ The diffusion model $p_{\theta}(x_{t},t|s)$ can then be trained to approximate the score function $\nabla_{x}\log q_{t}(x|s),$ which is used to simulate the noising process in reverse. We emphasize that ConforMix itself does not involve additional model training.
@@
is easy to compute. During inference, at each denoising timestep $t$, we construct an inexpensive estimate of the fully denoised state $\hat{x}_{0}$, which we plug into $y_{j}$ to approximate the potential of the current state. We then compute $\nabla_{x_{t}}y_{j}=(J_{x_{t}}\hat{x}_{0})^{T}\nabla_{\hat{x}_{0}}y_{j}$ and use this signal to guide $x_{t}$ towards $p(x|y_{j},s).$ By simultaneously denoising several such particles $x_{t}$ and periodically resampling them to maintain correct sampling, we can obtain approximate samples from the desired conditional distribution. For more details, we refer the reader to Algorithm S1 in the Supplement and to Algorithm 1 in [wu_practical_2024], which provides an asymptotic guarantee that the sampling approaches the true conditional distribution as the number of particles becomes large.
@@ 3.3 Sample reweighting and free energy estimation
To solve the challenge of estimating the partition functions $p(y_{j})$, we turn to a statistical tool that to the best of our knowledge has not previously been applied in the context of diffusion models. The multistate Bennett acceptance ratio (MBAR) free energy estimation algorithm combines sample information from multiple conditional distributions to jointly estimate the probabilities $\{p(y_{j})\},$ up to a (negligible) constant factor [shirts_statistically_2008, shirts_reweighting_2017]. In practice, this means that given sufficiently accurate samples from a series of overlapping distributions, such as the conditional distributions of ConforMix, we can reconstruct the unbiased model landscape in those areas. We discuss free energy estimation further in Section 4.4. The mathematical details of MBAR are deferred to the appendix.
@@ 4.1 ConforMixRMSD recovers experimentally observed conformations
We first implement ConforMix in Boltz-1 [wohlwend_boltz-1_2025], an open-source diffusion-based structure prediction model similar to AlphaFold 3. Importantly, Boltz was trained on the Protein Data Bank (PDB), and not with any other dynamics information. While the default Boltz sampling can generate multiple samples, they are generally highly concentrated and are more usefully thought of as multiple approximations of the same static structure than as samples from the true conformational distribution of the protein. To address Q1, we collect samples using ConforMixRMSD with Boltz (ConforMixRMSD-Boltz) as the underlying structure prediction model.
We examine performance of ConforMixRMSD-Boltz on proteins that exhibit different types of conformational changes: (a) domain motion, (b) transporter cycling, (c) cryptic pocket formation, and (d) fold switching. The set of 38 proteins that exhibit domain motions are the combination of proteins curated by Lewis et al. [lewis_scalable_2025] and in OC23 by Kalakoti et al. [kalakoti_afsample2_2025]. The set of 15 membrane transporters that exhibit inward-open and outward-open conformations were curated in TP16 by Kalaloti et al. [kalakoti_afsample2_2025]. The set of 31 proteins that exhibit cryptic pocket formation were curated by Lewis et al. [lewis_scalable_2025]. The set of 15 fold switchers proteins are a subset of those curated by Porter et al. [porter_extant_2018].
On these sets, Boltz alone typically recovers only a single conformation. Sampling from Boltz with ConforMixRMSD, however, reveals that the hidden energy landscape contains degrees of freedom that transition to alternative conformations seen in experiment and deposited in the PDB (Figure 2). As comparators, we implement three MSA-modification protocols in Boltz: AFCluster [wayment-steele_predicting_2024], AFSample2 [kalakoti_afsample2_2025], and CF-random [lee_large-scale_2025] (see details in Supplement). We also compare to default Boltz sampling with $1,000$ samples. On the domain motion, transport, and cryptic pocket sets, ConforMixRMSD recovers more alternate conformations than all baselines—as measured by RMSD and TM-scores to PDB structures (Tables 1, LABEL:supp-tab:metrics_auc and Figures 2, LABEL:supp-fig:rmsd-coverage, LABEL:supp-fig:tmscore, LABEL:supp-fig:rmsdgrid-domainmotion-LABEL:supp-fig:rmsdgrid-foldswitching)—demonstrating its power explore conformational space in a novel manner. MSA-based methods show stronger performance on fold switchers, suggesting such rearrangements are be better captured discrete input modulation, whereas ConforMixRMSD is better suited to continuous (e.g. domain, transporter, pocket) transitions.
@@
Worst-matched
reference
conformation
(harder task)
Default Boltz sampling 0.33 (±0.14) 0.13 (±0.17) 0.15 (±0.12) 0.13 (±0.17)
@@
Best-matched
reference
conformation
(easier task)
Default Boltz Sampling 0.94 (±0.07) 0.59 (±0.23) 0.94 (±0.08) 0.60 (±0.27)
@@
We also find that recovery of experimentally observed states from ConforMixRMSD-Boltz sampling is competitive with BioEmu [lewis_scalable_2025] (Figure LABEL:supp-rmsd_coverage_bioemu), a model trained to generate conformational ensembles, although we note that many of the evaluated proteins were held out of training for BioEmu but were likely present in the Boltz training set.
@@ 4.2 ConforMixRMSD samples conformational transitions of domain motion proteins
As an alternate metric, the “fill ratio" introduced by [kalakoti_afsample2_2025] measures how continuously samples cover the conformational path between reference states based on TM-scores. ConforMixRMSD sampling results in higher fill ratios than the other evaluated methods (Figure LABEL:supp-fig:fillratio). We also compute geometric quality metrics on ConforMix-generated samples and observe slightly elevated rates of outliers compared to default sampling, but most samples appear reasonable (Table LABEL:supp-tab:outliers).
@@ 4.3 ConforMixRMSD recovers heterogeneous states in multi-chain and RNA systems
We further evaluate ConforMixRMSD on select biomolecular systems known to exhibit interesting heterogeneity. First, we consider a multi-chain transporter protein, the SemiSWEET sugar transporter, which cycles through three distinct states to transport substrate across a membrane (Figure 4A) [latorraca_mechanism_2017]. The inward- and outward-open states enable substrate uptake and release, while an occluded state that is closed to both sides allows the substrate to pass through the transporter. Thus far, ML-based conformational sampling approaches have primarily been effective on single-chain proteins. Boltz by default predicts only the occluded state for this system. With ConforMixRMSD sampling, however, Boltz recovers all three states. PCA analysis and visual inspection of sampled structures reveal intermediate states between each known conformation, forming a conformational “trajectory".
While most of our testing focuses on proteins, ConforMix can be applied to any biomolecular system as long as it is supported by the underlying model. We demonstrate sampling of the hammerhead ribozyme (Figure 4B), a small catalytic RNA molecule that adopts multiple conformations. These conformations have not been deposited in the PDB, so instead we take molecular dynamics simulations from [lu2025hammerhead] as reference data (the grey structures in Figure 4B represent MD frames). ConforMixRMSD-Boltz visually recapitulates key aspects of the unzipping transition observed in the MD simulations. The sampled structures qualitatively capture the conformational trajectory, suggesting that ConforMix can extract meaningful conformational heterogeneity from RNA systems, though comprehensive evaluation across RNA systems would be needed to fully characterize performance.

### v2

@@ 1 Introduction
Biology depends on molecular flexibility. Proteins, RNA, DNA, and other components of biological systems adopt dynamic 3D conformations, which are key to function [8, 2]. Understanding and modeling these dynamics informs both basic biology research and applications to medicine, agriculture, and other areas, and is therefore a major aim of both experimental and computational scientists. The statistical distribution of the atomic configurations of a molecular system in equilibrium at a given temperature is determined by their potential energies. Specifically, a state with lower potential energy is exponentially more probable than a higher potential energy state, as determined by the Boltzmann distribution [32]. To sample this distribution, scientists have invested decades of effort in creating methods such as Monte Carlo samplers and molecular dynamics simulations (MD), but these methods are typically extremely slow and suffer from accuracy problems [32, 26, 33].
More recently, machine learning models have revolutionized biomolecular structure prediction, and efforts are underway to extend these advances to train models that do not merely predict static structures but generate diverse conformations from the true Boltzmann distribution [12, 1, 11, 36, 17]. As the field evolves, new questions arise about how best to take advantage of these models, access the information they contain, and target opportunities for further improvement. We ask the following question: Given a pretrained generative structure model, what strategy should be deployed at inference time to efficiently and accurately sample its conformational landscape?
In this work, we present an enhanced sampling algorithm, ConforMix, that can be applied to diffusion-based biomolecular structure prediction models such as the AlphaFold 3 family. ConforMix reveals hidden conformations and estimates free energies more efficiently than the default sampling from diffusion models. Our approach is based on the hypothesis that the probability distribution of generative models contains more information than is efficiently sampled. Importantly, even a generative model that perfectly sampled the true Boltzmann distribution—a major long-term goal in machine learning for biology—would, for many uses, still require enhanced sampling algorithms to fruitfully explore it. For instance, correct sampling of the tight-binding proteins barnase and barstar at standard state would generate an unbound structure with probability approximately $10^{-14}$ [34]. Directly drawing samples from the distribution would be an extremely inefficient way to estimate this probability, determine transition states along the binding pathway, or understand the effects of environmental factors in modulating this affinity [40, 31].
@@ 2 Related Work
Computational methods have been used to sample molecular conformations for decades. The well-known Metropolis-Hastings algorithm [24] for Monte Carlo sampling was first used to simulate the statistical mechanics of particles in a box. Elastic network models represent a protein as low-resolution beads connected by springs and recover the principal motions that would occur in this representation, which often are representative of motions of the actual system [21]. More sophisticated methods, such as coarse-grained or all-atom molecular dynamics simulation, model the physical interactions with progressively more detail and have been used to generate detailed probabilistic distributions for a number of proteins [19, 25].
More recently, machine learning models such as AlphaFold 3 [1] have revolutionized the problem of static protein structure prediction. AlphaFold 3 is the most prominent example of a class of diffusion-based structure prediction methods that excel at generating the single most likely protein structure for a given amino acid sequence. While AlphaFold 3 generates five samples per run, for well-ordered proteins these samples are often almost identical and are better understood as very similar predictions of the same static structure rather than a representation of conformational variety.
Using such models to model the distribution of conformations, however, is a task still in its infancy. The most well-studied approaches involve subsampling or otherwise perturbing the multiple sequence alignment (MSA) input [36, 28, 13, 15] or templates input [9, 5, 4] to AlphaFold and related models. These inputs specify, respectively, sequence co-evolutionary information and related 3D structures. Emphasizing different sequence or structural features in the input has empirically produced significant conformational diversity. Subsampling methods, however, suffer from fundamental limitations. First, subsampling necessarily involves reducing the information available to the model, which often produces poor-quality outputs. Second, molecular conformations are continuous, while subsampling of sequences or templates is a discrete operation. Third, the structure distribution accessed by subsampling methods is not well defined, and despite years of study there is no clear means of reconstructing actual probabilistic ensembles.
Beyond input-perturbation methods, other efforts have involved building new models to predict conformations [17, 7, 11]. Strategies include leveraging molecular dynamics simulations or experimental data, either during training or at runtime [20, 3, 35, 22, 16]. While these approaches have shown promise, they do not yet accurately reproduce the thermodynamic ensembles of arbitrary proteins.
@@ 3.1 ConforMix
Diffusion models learn probability distributions, and the standard sampling procedures, which we refer to as default sampling, are designed to draw i.i.d. samples from this distribution. To explore further and more efficiently than is possible with default sampling, we employ conditional sampling, which is widely used with diffusion models to guide sampling toward particular regions of the learned probability distribution [6, 10]. As we will see, conditional sampling offers clear benefits even without specialized prior knowledge about the protein of interest. While time-dependent classifier guidance and classifier-free guidance are commonly used in other diffusion contexts, these methods are not well suited to our needs because they require either additional training or special architectural choices, which are impractical for our needs.
Instead, to devise a purely inference-time method, we turn to Twisted Diffusion Sampling [38], a form of sequential Monte Carlo (SMC), to guide conditional generation. SMC has advantages over more typical guidance methods, including (a) requiring only a time-independent guidance function, which enables us to easily describe conditions as a function of the fully denoised state of the system, and (b) providing an asymptotic guarantee of sampling the true target conditional distribution in the limit of many particles. Twisted diffusion sampling works in conjunction with an underlying diffusion model that learns to sample a target data distribution $q(x_{0}|s).$ Diffusion models define a forward noising process, operating over a sequence of $T$ steps, that constructs an extended distribution $q(x_{T:0}|s)=q(x_{T}|x_{T_{1}})\cdots q(x_{1}|x_{0})q(x_{0}|s).$ The diffusion model $p_{\theta}(x_{t},t|s)$ can then be trained to approximate the score function $\nabla_{x}\log q_{t}(x|s),$ which is used to simulate the noising process in reverse. We emphasize that ConforMix itself does not involve additional model training.
@@
is easy to compute. During inference, at each denoising timestep $t$, we construct an inexpensive estimate of the fully denoised state $\hat{x}_{0}$, which we plug into $y_{j}$ to approximate the potential of the current state. We then compute $\nabla_{x_{t}}y_{j}=(J_{x_{t}}\hat{x}_{0})^{T}\nabla_{\hat{x}_{0}}y_{j}$ and use this signal to guide $x_{t}$ towards $p(x|y_{j},s).$ By simultaneously denoising several such particles $x_{t}$ and periodically resampling them to maintain correct sampling, we can obtain approximate samples from the desired conditional distribution. For more details, we refer the reader to Algorithm S1 in the Supplement and to Algorithm 1 in [38], which provides an asymptotic guarantee that the sampling approaches the true conditional distribution as the number of particles becomes large.
@@ 3.3 Sample reweighting and free energy estimation
To solve the challenge of estimating the partition functions $p(y_{j})$, we turn to a statistical tool that to the best of our knowledge has not previously been applied in the context of diffusion models. The multistate Bennett acceptance ratio (MBAR) free energy estimation algorithm combines sample information from multiple conditional distributions to jointly estimate the probabilities $\{p(y_{j})\},$ up to a (negligible) constant factor [29, 30]. In practice, this means that given sufficiently accurate samples from a series of overlapping distributions, such as the conditional distributions of ConforMix, we can reconstruct the unbiased model landscape in those areas. We discuss free energy estimation further in Section 4.4. The mathematical details of MBAR are deferred to the appendix.
@@ 4.1 ConforMixRMSD recovers experimentally observed conformations
We first implement ConforMix in Boltz-1 [37], an open-source diffusion-based structure prediction model similar to AlphaFold 3. Importantly, Boltz was trained on the Protein Data Bank (PDB), and not with any other dynamics information. While the default Boltz sampling can generate multiple samples, they are generally highly concentrated and are more usefully thought of as multiple approximations of the same static structure than as samples from the true conformational distribution of the protein. To address Q1, we collect samples using ConforMixRMSD with Boltz (ConforMixRMSD-Boltz) as the underlying structure prediction model.
We examine performance of ConforMixRMSD-Boltz on proteins that exhibit different types of conformational changes: (a) domain motion, (b) transporter cycling, (c) cryptic pocket formation, and (d) fold switching. The set of 38 proteins that exhibit domain motions are the combination of proteins curated by Lewis et al. [17] and in OC23 by Kalakoti et al. [13]. The set of 15 membrane transporters that exhibit inward-open and outward-open conformations were curated in TP16 by Xie & Huang [39]. The set of 31 proteins that exhibit cryptic pocket formation were curated by Lewis et al. [17]. The set of 15 fold switchers proteins are a subset of those curated by Porter et al. [27].
On these sets, Boltz alone typically recovers only a single conformation. Sampling from Boltz with ConforMixRMSD, however, reveals that the hidden energy landscape contains degrees of freedom that transition to alternative conformations seen in experiment and deposited in the PDB (Figure 2). As comparators, we implement three MSA-modification protocols in Boltz: AFCluster [36], AFSample2 [13], and CF-random [15] (see details in Supplement). We also compare to default Boltz sampling with $1,000$ samples. On the domain motion, transport, and cryptic pocket sets, ConforMixRMSD recovers more alternate conformations than all baselines—as measured by RMSD and TM-scores to PDB structures (Tables 1, LABEL:supp-tab:metrics_auc and Figures 2, LABEL:supp-fig:rmsd-coverage, LABEL:supp-fig:tmscore, LABEL:supp-fig:rmsdgrid-domainmotion-LABEL:supp-fig:rmsdgrid-foldswitching)—demonstrating its power explore conformational space in a novel manner. MSA-based methods show stronger performance on fold switchers, suggesting such rearrangements are be better captured discrete input modulation, whereas ConforMixRMSD is better suited to continuous (e.g. domain, transporter, pocket) transitions.
@@
Worst-matched reference conformation (harder task) Default Boltz sampling 0.33 (±0.14) 0.13 (±0.17) 0.15 (±0.12) 0.13 (±0.17)
@@
Best-matched reference conformation (easier task) Default Boltz Sampling 0.94 (±0.07) 0.59 (±0.23) 0.94 (±0.08) 0.60 (±0.27)
@@
We also find that recovery of experimentally observed states from ConforMixRMSD-Boltz sampling is competitive with BioEmu [17] (Figure LABEL:supp-rmsd_coverage_bioemu), a model trained to generate conformational ensembles, although we note that many of the evaluated proteins were held out of training for BioEmu but were likely present in the Boltz training set.
@@ 4.2 ConforMixRMSD samples conformational transitions of domain motion proteins
As an alternate metric, the “fill ratio" introduced by [13] measures how continuously samples cover the conformational path between reference states based on TM-scores. ConforMixRMSD sampling results in higher fill ratios than the other evaluated methods (Figure LABEL:supp-fig:fillratio). We also compute geometric quality metrics on ConforMix-generated samples and observe slightly elevated rates of outliers compared to default sampling, but most samples appear reasonable (Table LABEL:supp-tab:outliers).
@@ 4.3 ConforMixRMSD recovers heterogeneous states in multi-chain and RNA systems
We further evaluate ConforMixRMSD on select biomolecular systems known to exhibit interesting heterogeneity. First, we consider a multi-chain transporter protein, the SemiSWEET sugar transporter, which cycles through three distinct states to transport substrate across a membrane (Figure 4A) [14]. The inward- and outward-open states enable substrate uptake and release, while an occluded state that is closed to both sides allows the substrate to pass through the transporter. Thus far, ML-based conformational sampling approaches have primarily been effective on single-chain proteins. Boltz by default predicts only the occluded state for this system. With ConforMixRMSD sampling, however, Boltz recovers all three states. PCA analysis and visual inspection of sampled structures reveal intermediate states between each known conformation, forming a conformational “trajectory".
While most of our testing focuses on proteins, ConforMix can be applied to any biomolecular system as long as it is supported by the underlying model. We demonstrate sampling of the hammerhead ribozyme (Figure 4B), a small catalytic RNA molecule that adopts multiple conformations. These conformations have not been deposited in the PDB, so instead we take molecular dynamics simulations from [23] as reference data (the grey structures in Figure 4B represent MD frames). ConforMixRMSD-Boltz visually recapitulates key aspects of the unzipping transition observed in the MD simulations. The sampled structures qualitatively capture the conformational trajectory, suggesting that ConforMix can extract meaningful conformational heterogeneity from RNA systems, though comprehensive evaluation across RNA systems would be needed to fully characterize performance.